# on the Radon-Nikodym Derivative

### v1

@@ Abstract
Rigorous statements and formal proofs are presented for both foundational and advanced folklore theorems on the Radon-Nikodym derivative. The cases of product and marginal measures are carefully considered; and the hypothesis under which the statements hold are rigorously enumerated.
@@ I Introduction
In mathematics, folklore theorems refer to results that are widely accepted and frequently utilized by experts but are often not formally proven or explicitly documented. In game theory for example, the original Folk theorem earned its name because, although it was widely recognized among game theorists during the 1950s, it remains unpublished and without attribution to particular authors. See for instance [1] and [2]. Folklore theorems populate all areas in mathematics. In information theory, a large set of folklore theorems involve the Radon-Nikodym derivative (RND), first introduced by Radon [3]; and later generalized by Nikodym [4]. The existence of such folklore theorems in this area arises in part from the fact that all Shannon’s information measures can be defined in terms of the RND. Interestingly, most properties of the RND are often presented as comments in most textbooks on measure theory and probability theory, c.f, [5, 6, 7, 8, 9, 10, 11]
Claude Shannon did not use the RND in his foundational papers [12, 13] to define entropy and mutual information. Instead, Shannon opted for restricting his publications to the case in which measures either possess a probability mass function or a probability density function, which are both instances of RNDs. This choice significantly influenced the presentation of most subsequent results in information theory and established the style in which classical textbooks were written [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Nonetheless, the RND has been increasingly adopted in modern textbooks [31] and in the definition of new information measures, e.g., lautum information [32], to privilege a unified presentation. That is, independently of the measure used as a reference, e.g., the counting measure, the Lesbegue measure, etc. Some recent results whose presentation relies on the RND are for instance [33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. In the light of the central role of the RND in information theory, this paper presents rigorous statements and formal proofs for the most common folklore theorems on the RND. These include the change of measure theorem, the proportional-measures theorem, the chain rule, the multiplicative inverse, linearity and continuity theorems, as well as the product-measure theorem. Less common theorems such as the unit measure and two Bayes-like rules of the RND are also formally proved. This is the first time such material appears in peer-reviewed literature. Finally, it is important to highlight that particular attention has been put on rigorously stating the conditions under which these theorems hold, hopefully providing a valuable reference for researchers and students in information theory.
@@ II Preliminaries
Theorem 1 (Radon-Nikodym theorem, [5, Theorem 2.2.1])
Let $P$ and $Q$ be two measures on a given measurable space $\left(\Omega,\mathscr{F}\right)$, such that $Q$ is $\sigma$-finite and ${P}\ll{Q}$. Then, there exists a nonnegative Borel measurable function $g:~{}\Omega~{}\to~{}\mathbb{R}$ such that for all $\mathcal{A}\in\mathscr{F}$,
@@
The function $g$ in (1) is often referred to as the Radom-Nikodym derivative of $P$ with respect to $Q$; and is also written as $\frac{\mathrm{d}P}{\mathrm{d}Q}$, such that $g(x)~{}=~{}\frac{\mathrm{d}P}{\mathrm{d}Q}(x).$
The Radon-Nikodym theorem is the foundational tool from which many folklore theorems in information theory originate. Some of these folklore theorems are thoroughly studied in the following sections.
@@ III Basic Folklore Theorems
Theorem 2 (Change of Measure)
Let $P$ and $Q$ be two measures on the measurable space $\left(\Omega,\mathscr{F}\right)$ with ${P}\ll{Q}$; and $Q$ a $\sigma$-finite measure. Let $f:\Omega\to\mathds{R}$ be a Borel measurable function such that the integral $\int_{\Omega}f(x)\mathrm{d}P(x)$ exists. Then, for all $\mathcal{A}\in\mathscr{F}$,
@@
[…]
@@
$=$ $\sum_{i=1}^{n}a_{i}\int_{\mathcal{B}_{i}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)$ (4)
$=$ $\sum_{i=1}^{n}a_{i}P(\mathcal{B}_{i}),$ (5)
where the equality in (LABEL:EqMartinezPabon102) follows from the linearity of the integral [5, Theorem 1.6.3]; and the equality in (LABEL:EqMartinezPabon105) follows from Theorem 1. On the other hand, for all $\mathcal{A}\in\mathscr{F}$
$\displaystyle\int_{\mathcal{A}}f(x)\mathrm{d}P(x)$ $=$ $\displaystyle\int_{\mathcal{A}}\displaystyle\sum_{i=1}^{n}a_{i}\mathds{1}_{\mathcal{A}_{i}(x)}\mathrm{d}P(x)$ (6)
$=$ $\displaystyle\sum_{i=1}^{n}\int_{\mathcal{A}}a_{i}\mathds{1}_{\mathcal{A}_{i}(x)}\mathrm{d}P(x)$ (7)
$=$ $\sum_{i=1}^{n}a_{i}\int_{\mathcal{B}_{i}}\mathrm{d}P(x)=\sum_{i=1}^{n}a_{i}P(\mathcal{B}_{i}),$ (8)
where the equality in (7) follows from the linearity of the integral [5, Theorem 1.6.3]. Hence, from Theorem 1, and equalities (LABEL:EqMartinezPabon105) and (8), it follows that when $f$ is a simple function, the equality in (2) holds. This concludes the first part of the proof.
The second part of the proof proceeds by considering the following observations: $\left(a\right)$ simple functions form a dense subset of the space of Borel measurable functions [5, Theorem 1.5.5(b)]; and $\left(b\right)$ the integral is a continuous map from that space [5, Theorem 1.6.2]. Hence, from $\left(a\right)$ and $\left(b\right)$, it follows that (2) also holds for any Borel measurable function $f$. This completes the proof. ∎
Another reputed folklore theorem, which is often referred to as the “proportional measures” theorem, establishes the explicit forms of the Radon-Nikodym derivatives between two measures, in which one is proportional to the other.
Theorem 3 (Proportional Measures)
@@
with $c>0$. Then, for all $x\in\Omega$
$\frac{\mathrm{d}P}{\mathrm{d}Q}(x)$ =$~{}\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{1}{c},\mbox{ and }\frac{\mathrm{d}Q}{\mathrm{d}P}(x)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}c.$ (10)
@@
First, note that (9) implies that the measures $P$ and $Q$ are mutually absolutely continuous. Hence, from Theorem 1, it follows that for all $\mathcal{A}\in\mathscr{F}$,
$P\left(\mathcal{A}\right)=\int_{\mathcal{A}}\mathrm{d}P(x)=\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\;\mathrm{d}Q(x).$ (11)
On the other hand, the equality (9) also implies
$P\left(\mathcal{A}\right)=\frac{1}{c}Q(\mathcal{A})=\int_{\mathcal{A}}\frac{1}{c}\mathrm{d}Q(x).$ (12)
Hence, it follows directly from Theorem 1 that the Radon-Nikodym derivative $\frac{\mathrm{d}P}{\mathrm{d}Q}$ is unique almost surely with respect to $Q$. Thus, $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$$\frac{1}{c}$. Using similar arguments and the fact that $P$ and $Q$ are mutually absolutely continuous, it is verified that $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$$c$. ∎
In the case in which $c=1$ in (10), measures $P$ and $Q$ are identical, thus, $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$ $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$ $1$
The following folklore theorem is often referred to as the “chain rule”.
Theorem 4 (Chain Rule)
Let $P$, $Q$, and $R$ be three measures on the measurable space $\left(\Omega,\mathscr{F}\right)$ such that ${P}\ll{Q}$; ${Q}\ll{R}$; and $Q$ and $R$ are $\sigma$-finite measures. Then,
@@
From the assumptions of the theorem, it follows that for all $\mathcal{A}\in\mathscr{F}$,
$P(\mathcal{A})$ $=$ $\int_{\mathcal{A}}\mathrm{d}P(x)=\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)$ (14)
$=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\frac{\mathrm{d}Q}{\mathrm{d}R}(x)\mathrm{d}R(x)$ (15)
$=$ $\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}R}(x)\mathrm{d}R(x),$ (16)
where the second equality in (14) follows from Theorem 1; and the equality in (15) follows from Theorem 2 . The equality in (16) holds from Theorem 1 and by noticing that $P~{}\ll~{}R$. Therefore, the equalities in (15) and (16) together with Theorem 1 imply (13), which completes the proof. ∎
The following folklore theorem shows the connection between the Radon-Nikodym derivative and its multiplicative inverse.
Theorem 5 (Multiplicative Inverse)
Let $P$ and $Q$ be two mutually absolutely continuous measures on the measurable space $\left(\Omega,\mathscr{F}\right)$; and assume that for all $x\in\Omega$, $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)>0$. Then,
Proof:
From Theorem 4, it follows that
$\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{\mathrm{d}Q}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}1,$ (18)
where the last equality follows from Theorem 3, with $c~{}=~{}1$. This completes the proof. ∎
The subsequent folklore theorem establishes the linearity of the Radon-Nikodym derivative.
Theorem 6 (Linearity)
Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$ and let also $Q_{1},Q_{2},\ldots,Q_{n}$ be finite measures on $\left(\Omega,\mathscr{F}\right)$ absolutely continuous with respect to $P$. Let $c_{1},c_{2},\ldots,c_{n}$ be positive reals; and let $S$ be a finite measure on $\left(\Omega,\mathscr{F}\right)$ such that for all $\mathcal{A}\in\mathscr{F}$, $S(\mathcal{A})=\sum_{t=1}^{n}c_{t}Q_{t}(\mathcal{A})$. Then,
$\frac{\mathrm{d}S}{\mathrm{d}P}(x)$ =$\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$ $\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x).$ (19)
Proof:
The proof starts by noticing that, from the assumptions of the theorem, it holds that ${S}\ll{P}$. Hence, for all $\mathcal{A}\in\mathscr{F}$, it holds that
$\int_{\mathcal{A}}\frac{\mathrm{d}S}{\mathrm{d}P}(x)\mathrm{d}P(x)=\int_{\mathcal{A}}\mathrm{d}S(x)=\sum_{t=1}^{n}c_{t}Q_{t}(\mathcal{A})$ (20)
$=$ $\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\mathrm{d}Q_{t}(x)=\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x)$ (21)
$=$ $\int_{\mathcal{A}}\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x),$ (22)
where the first equality in (20) and the last equality in (21) follow from Theorem 2; and the equality (22) follows from the additivity property of the integral [5, Corollary 1.6.4]. The proof ends by using Theorem 1, which implies the equality in (19) from (22). ∎
@@
Theorem 7 (Continuity)
Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$, and let $Q_{1},Q_{2},\cdots$ be an infinite sequence of $\sigma$-finite measures on $\left(\Omega,\mathscr{F}\right)$, converging to a measure $Q$. Suppose that for all $n\in\mathbb{N}$, ${Q_{n}}\ll{P}$. Then, ${Q}\ll{P}$ and
@@
From the assumptions of the theorem, for all $\mathcal{A}\in\mathscr{F}$, it holds that
$Q(\mathcal{A})$ $=$ $\lim_{n\rightarrow\infty}Q_{n}(\mathcal{A})$ (24)
$=$ $\lim_{n\rightarrow\infty}\int_{\mathcal{A}}\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(x)\mathrm{d}P(x)$ (25)
$=$ $\int_{\mathcal{A}}\lim_{n\to\infty}\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(x)\mathrm{d}P(x),$ (26)
where the equality in (25) follows from Theorem 2 ; and the equality in (26) follows from [5, Theorem 1.6.2]. The equality in (26) implies that ${Q}\ll{P}$. Hence, for all $\mathcal{A}\in\mathscr{F}$, it holds that
@@
Therefore, the equalities in (26) and (27) jointly with Theorem 1 imply equation (23), which completes the proof. ∎
The ensuing folklore theorem establishes the relation between the Radon-Nikodym derivative of a product measure with respect to its component measures.
Theorem 8 (Product of Measures)
For all $i\in\{1,2\}$, let $P_{i}$ and $Q_{i}$ be a finite and a $\sigma$-finite measure on $\left(\Omega_{i},\mathscr{F}_{i}\right)$, respectively; with ${P_{i}}\ll{Q_{i}}$. Let also $P_{1}P_{2}$ and $Q_{1}Q_{2}$ be the product measures on $\left(\Omega_{1}\times\Omega_{2},\mathscr{F}_{1}\times\mathscr{F}_{2}\right)$ formed by $P_{1}$ and $P_{2}$; and $Q_{1}$ and $Q_{2}$, respectively. Then,
@@
From the assumptions of the theorem, for all $\mathcal{A}\in\left(\Omega_{1}\times\Omega_{2}\right)$,
$P_{1}P_{2}\left(\mathcal{A}\right)$ $=$ $\int_{\mathcal{A}}\mathrm{d}P_{1}P_{2}\left(x_{1},x_{2}\right)$ (29)
$=$ $\int\int_{\mathcal{A}_{x_{2}}}\mathrm{d}P_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (30)
$=$ $\int\int_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (31)
$=$ $\int\int_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\dfrac{\mathrm{d}P_{2}\left(x_{2}\right)}{\mathrm{d}Q_{2}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}Q_{2}\left(x_{2}\right)$ (32)
@@
where $\mathcal{A}_{x_{2}}$ is the section of the set $\mathcal{A}$ determined by $x_{2}$, namely, $\mathcal{A}_{x_{2}}\triangleq\left\{x_{1}\in\Omega_{1}:(x_{1},x_{2})\in\mathcal{A}\right\}$; the equality in (29) arises from the definition of $P_{1}P_{2}$ as the product of $P_{1}$ and $P_{2}$; the equality in (31) is a direct consequence of Theorem 2; the equality in (32) follows from Theorem 1; and finally, the equality in (33) is due to the construction of $Q_{1}Q_{2}$ as the product measure of $Q_{1}$ and $Q_{2}$.
The proof follows by observing that from the equality in (33), it holds that ${P_{1}P_{2}}\ll{Q_{1}Q_{2}}$. Thus, for all $\mathcal{A}~{}\in~{}\mathscr{F}_{1}~{}\times~{}\mathscr{F}_{2}$,
@@
The equalities in (33) and (34), together with Theorem 1, imply the equality in (28), which completes the proof. ∎

### v2

@@ Abstract
In this paper, rigorous statements and formal proofs are presented for both foundational and advanced folklore theorems on the Radon-Nikodym derivative. The cases of conditional and marginal probability measures are carefully considered, which leads to an identity involving the sum of mutual and lautum information suggesting a new interpretation for such a sum.
@@ I Introduction
In mathematics, folklore theorems refer to results that are widely accepted and frequently utilized by experts but are often not formally proven or explicitly documented. In game theory for example, the original Folk theorem earned its name because, although it was widely recognized among game theorists during the 1950s, it remains unpublished and without attribution to particular authors. See for instance [1] and [2]. Folklore theorems populate all areas in mathematics. In information theory, a large set of folklore theorems involve the Radon-Nikodym derivative (RND), first introduced by Radon [3]; and later generalized by Nikodym [4]. The existence of such folklore theorems in this area arises in part from the fact that all Shannon’s information measures can be defined in terms of the RND. Interestingly, most properties of the RND are often presented without proof in most textbooks on measure theory and probability theory, c.f, [5, 6, 7, 8, 9, 10, 11], with the exception of some properties presented in [12].
Claude Shannon did not use the RND in his foundational papers [13, 14] to define entropy and mutual information. Instead, Shannon opted for restricting his work to the case in which measures either possess a probability mass function or a probability density function, which are both instances of RNDs. This choice significantly influenced the presentation of most subsequent results in information theory and established the style in which classical textbooks were written [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Nonetheless, the RND has been increasingly adopted in modern textbooks [32] and in the definition of new information measures, e.g., lautum information [33], to privilege a unified presentation. That is, independently of the measure used as a reference, e.g., the counting measure, the Lesbegue measure, etc. Some recent results whose presentation relies on the RND are for instance [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. At the light of this observation, this paper presents rigorous statements and alternative proofs for the most common folklore theorems on the RND. These include the change of measure theorem, the proportional-measures theorem, the chain rule, the multiplicative inverse, linearity and continuity theorems, as well as the product-measure theorem. Less common theorems such as the unit measure and two Bayes-like rules of the RND are also formally proved for the first time in peer-reviewed literature. Finally, these novel theorems are used to prove a key identity involving the sum of mutual and lautum information, which is often observed in statistical learning [46, 42]. Hopefully, these contributions would provide a valuable reference for researchers and students in information theory.
@@ II Preliminaries
Theorem 1 (Radon-Nikodym theorem, [5, Theorem 2.2.1]).
Let $P$ be a signed measure and $Q$ be a $\sigma$-finite measure both on a given measurable space $\left(\Omega,\mathscr{F}\right)$, with ${P}\ll{Q}$. Then, there exists a Borel measurable function $g:\leavevmode\nobreak\ \Omega\leavevmode\nobreak\ \to\leavevmode\nobreak\ \mathbb{R}$ such that for all $\mathcal{A}\in\mathscr{F}$,
@@
The function $g$ in (1) is often referred to as the Radom-Nikodym derivative of $P$ with respect to $Q$; and is also written as $\frac{\mathrm{d}P}{\mathrm{d}Q}$, such that $g(x)\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\mathrm{d}P}{\mathrm{d}Q}(x).$
The Radon-Nikodym derivative exhibits a number of properties from which many foundational results originate. Some of these properties are certainly folklore theorems and thus, the purpose of the following section is to provide formal proofs. Other properties are formally proved in literature and thus, such properties are simply stated for the sake of completeness. In some cases, alternative proofs are presented, for privileging simplicity.
@@ III Basic Folklore Theorems
Theorem 2 (Change of Measure).
Let $P$ be a signed measure and $Q$ be a $\sigma$-finite measure both on a given measurable space $\left(\Omega,\mathscr{F}\right)$, with ${P}\ll{Q}$. Let $f:\Omega\to\mathbb{R}$ be a Borel measurable function such that the integral $\int_{\Omega}f(x)\mathrm{d}P(x)$ exists. Then, for all $\mathcal{A}\in\mathscr{F}$,
@@
A proof of this theorem is presented in [12, Proposition 3.9]. For the sake of completeness, an alternative proof is presented hereunder, using the previously described notation.
@@
Assume that $f$ is a simple function, i.e. for all $x\in\Omega$, $f(x)=\sum_{i=1}^{m}a_{i}\mathds{1}_{\mathcal{A}_{i}}(x),$ for some finite $m\in\mathbb{N}$, disjoint sets $\mathcal{A}_{1},\ldots,\mathcal{A}_{m}$ in $\mathscr{F}$, and reals $a_{1},\ldots,a_{m}$. For all $\mathcal{A}\in\mathscr{F}$ and $i\in\{1,\ldots,m\}$, let $\mathcal{B}_{i}=\mathcal{A}\cap\mathcal{A}_{i}$. Then,
@@
$=$ $\sum_{i=1}^{n}a_{i}\int_{\mathcal{B}_{i}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\mathrm{d}Q(x)=\sum_{i=1}^{n}a_{i}P(\mathcal{B}_{i}),$ (4)
where the equality in (4) follows Theorem 1 and the linearity of the integral [5, Theorem 1.6.3]. On the other hand, for all $\mathcal{A}\in\mathscr{F}$,
$\int_{\mathcal{A}}f(x)\,\mathrm{d}P(x)$ $=$ $\sum_{i=1}^{m}a_{i}\int_{\mathcal{B}_{i}}\mathrm{d}P(x)\,=\,\sum_{i=1}^{m}a_{i}P(\mathcal{B}_{i}),$ (5)
where the equality in (5) follows from the linearity of the integral [5, Theorem 1.6.3]. Hence, from Theorem 1 when $f$ is simple, the desired equality in (2) holds immediately.
The proof proceeds by observing that: (a) simple functions are dense in the space of Borel measurable functions [5, Theorem 1.5.5(b)], and (b) the integral is a continuous map on that space [5, Theorem 1.6.2]. From (a) and (b), it follows that (2) also holds for any Borel measurable function $f$. This completes the proof. ∎
Another reputed folklore theorem, which is often referred to as the “proportional measures” theorem, establishes the explicit form of the Radon-Nikodym derivative between two measures, in which one is proportional to the other.
Theorem 3 (Proportional Measures).
@@
with $c>0$. Then,
$\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{1}{c},\quad\text{and}\quad\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}c.$ (7)
@@
Note that from (6), $P$ and $Q$ are mutually absolutely continuous. Moreover, from Theorem 1, for all $\mathcal{A}\in\mathscr{F}$,
$P(\mathcal{A})=\int_{\mathcal{A}}\mathrm{d}P(x)\;=\;\int_{\mathcal{A}}\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\,\mathrm{d}Q(x),$ (8)
and from (6), it follows that $P(\mathcal{A})=\frac{1}{c}\,Q(\mathcal{A})\;=\;\int_{\mathcal{A}}\frac{1}{c}\,\mathrm{d}Q(x)$. Hence, $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}\frac{1}{c}$ follows directly from Theorem 1. A similar argument shows that $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}c$. ∎
In the special case $c=1$ in (7), $P$ and $Q$ coincide, yielding $\frac{\mathrm{d}P}{\mathrm{d}Q}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{Q}$ $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}$ $1$
The following folklore theorem is often referred to as the “chain rule” and a proof appears in [12, Proposition 3.9].
Theorem 4 (Chain Rule).
Let $P$ be a signed measure and let $Q$ and $R$ be two $\sigma$-finite measures, all on the same measurable space $\left(\Omega,\mathscr{F}\right)$, with ${P}\ll{Q}$ and ${Q}\ll{R}$. Then,
@@
The following folklore theorem is a consequence of Theorem 4 and shows the connection between the Radon-Nikodym derivative and its multiplicative inverse. This result also appears in [12, Corollary 3.10].
Theorem 5 (Multiplicative Inverse).
Let $P$ and $Q$ be two mutually absolutely continuous $\sigma$-finite measures on the measurable space $\left(\Omega,\mathscr{F}\right)$; and assume that for all $x\in\Omega$, $\frac{\mathrm{d}Q}{\mathrm{d}P}(x)>0$. Then,
The subsequent folklore theorem establishes the linearity of the Radon-Nikodym derivative. A similar statement is presented in [12, Proposition 3.11] but left without proof.
Theorem 6 (Linearity).
Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$ and let also $Q_{1},Q_{2},\ldots,Q_{n}$ be signed measures on $\left(\Omega,\mathscr{F}\right)$ absolutely continuous with respect to $P$. Let $c_{1},c_{2},\ldots,c_{n}$ be positive reals; and let $S$ be a signed measure on $\left(\Omega,\mathscr{F}\right)$ such that for all $\mathcal{A}\in\mathscr{F}$, $S(\mathcal{A})=\sum_{t=1}^{n}c_{t}Q_{t}(\mathcal{A})$. Then,
$\frac{\mathrm{d}S}{\mathrm{d}P}(x)\displaystyle\mathop{\stackrel{{\scriptstyle a.s.}}{{=}}}_{P}\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x).$ (11)
@@
From the assumptions of the theorem, it holds that ${S}\ll{P}$, thus for all $\mathcal{A}\in\mathscr{F}$,
$\int_{\mathcal{A}}$ $\frac{\mathrm{d}S}{\mathrm{d}P}(x)\mathrm{d}P(x)=\int_{\mathcal{A}}\mathrm{d}S(x)=\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\mathrm{d}Q_{t}(x)$ (12)
$=$ $\sum_{t=1}^{n}\int_{\mathcal{A}}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x)=\int_{\mathcal{A}}\sum_{t=1}^{n}c_{t}\frac{\mathrm{d}Q_{t}}{\mathrm{d}P}(x)\mathrm{d}P(x),$ (13)
where the first equality in (12) and the first equality in (13) follow from Theorem 2; and the last equality in (13) follows from the additivity property of the integral [5, Corollary 1.6.4]. The proof ends by using Theorem 1 yielding (11). ∎
@@
Theorem 7 (Continuity).
Let $P$ be a $\sigma$-finite measure on $\left(\Omega,\mathscr{F}\right)$, and let $Q_{1},Q_{2},\cdots$ be an infinite sequence of signed measures on $\left(\Omega,\mathscr{F}\right)$, converging to a signed measure $Q$. Suppose that for all $n\in\mathbb{N}$, ${Q_{n}}\ll{P}$. Then, ${Q}\ll{P}$ and
@@
From the assumptions of the theorem and Theorem 2, for all $\mathcal{A}\in\mathscr{F}$, it holds that
$Q(\mathcal{A})$ $=$ $\lim_{n\to\infty}Q_{n}(\mathcal{A})\;=\;\int_{\mathcal{A}}\lim_{n\to\infty}\frac{\mathrm{d}Q_{n}}{\mathrm{d}P}(x)\,\mathrm{d}P(x),$ (15)
where the last equality in (15) follows from Theorem 2 and [5, Theorem 1.6.2]. Hence $Q\ll P$ and for all $\mathcal{A}\in\mathscr{F}$, it holds that
@@
Therefore, the equality in (16) and Theorem 1 yields (14). ∎
The ensuing folklore theorem relates the Radon–Nikodym derivative of a product measure to those of its component measures. This result also appears in [12, Exercise 3.12].
Theorem 8 (Product of Measures).
For all $i\in\{1,2\}$, let $P_{i}$ and $Q_{i}$ be a signed and a $\sigma$-finite measure on $\left(\Omega_{i},\mathscr{F}_{i}\right)$, respectively; with ${P_{i}}\ll{Q_{i}}$. Let also $P_{1}P_{2}$ and $Q_{1}Q_{2}$ be the product measures on $\left(\Omega_{1}\times\Omega_{2},\mathscr{F}_{1}\times\mathscr{F}_{2}\right)$ formed by $P_{1}$ and $P_{2}$; and $Q_{1}$ and $Q_{2}$, respectively. Then,
@@
By definition of product measures, for all $\mathcal{A}\subseteq\Omega_{1}\times\Omega_{2}$,
$P_{1}P_{2}\left(\mathcal{A}\right)$ $=$ $\int_{\mathcal{A}}\mathrm{d}P_{1}P_{2}\left(x_{1},x_{2}\right)=\iint_{\mathcal{A}_{x_{2}}}\mathrm{d}P_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (18)
$=$ $\iint_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}P_{2}\left(x_{2}\right)$ (19)
$=$ $\iint_{\mathcal{A}_{x_{2}}}\dfrac{\mathrm{d}P_{1}\left(x_{1}\right)}{\mathrm{d}Q_{1}}\dfrac{\mathrm{d}P_{2}\left(x_{2}\right)}{\mathrm{d}Q_{2}}\mathrm{d}Q_{1}\left(x_{1}\right)\mathrm{d}Q_{2}\left(x_{2}\right)$ (20)
@@
where $\mathcal{A}_{x_{2}}$ is the section of the set $\mathcal{A}$ determined by $x_{2}$, namely, $\mathcal{A}_{x_{2}}=\{x_{1}:(x_{1},x_{2})\in\mathcal{A}\}$; the equality in (18) arises from the definition of $P_{1}P_{2}$ as the product of $P_{1}$ and $P_{2}$; the equalities in (19) and (20) are direct consequences of Theorem 2; and finally, the equality in (21) is due to the construction of $Q_{1}Q_{2}$ as the product measure of $Q_{1}$ and $Q_{2}$. From the equality in (21), it holds that ${P_{1}P_{2}}\ll{Q_{1}Q_{2}}$. Thus, for all $\mathcal{A}\leavevmode\nobreak\ \in\leavevmode\nobreak\ \mathscr{F}_{1}\leavevmode\nobreak\ \times\leavevmode\nobreak\ \mathscr{F}_{2}$,
@@
From Theorem 1, the equality in (17) follows, completing the proof. ∎